OpenStudy (trwatkins1):

Macy measured the mass of a 69.3 g mouse at 70.0 g. Eli measured the mass of a 3867 kg elephant at 3900 kilograms. Macy's error was 70-69.3 = 0.7 g-- about half the mass of a paperclip. A. What was Eli's error? 3900-3867=33 kg B. Name an object of this approximate mass. about 9 gallons of water C. Accuracy can be defined in terms of percent error. Who was more accurate, Macy or Eli? Justify your answer with mathematics. 0.7 g more than 33kg

OpenStudy (trwatkins1):

@hartnn can you check my answers real quick please

hartnn (hartnn):

A,B correct

OpenStudy (trwatkins1):

:)

OpenStudy (trwatkins1):

how about c

hartnn (hartnn):

% error = (error/actual)*100

hartnn (hartnn):

for macy, error% = (0.7/69.3)*100 = ... for Eli,it error% = (33/3867)*100 = ..

hartnn (hartnn):

macy did 1% error, Eli di 0.8 % error macy did more erro :)

OpenStudy (trwatkins1):

macy

hartnn (hartnn):

means Eli is more accurate